Please help me out!
Let f(x)=x to the second power -16 and g(x)=4-x.
Find (fg)(-3).
(fg)(-3)=

Answer by Theo(13342)

Let f(x)=x to the second power -16 and g(x)=4-x.
Find (fg)(-3).

f(x) = x^2 - 16
g(x) = 4-x

fg(x) = f(x) * g(x)

that becomes fg(x) = (x^2 - 16) * (4-x)

fg(-3) becomes ((-3)^2 - 16) * (4 - (-3))

you simply replace x with (-3) in the equation.

solve that to get fg(-3) = (-7) * (7) = (-49)